Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below: Sales (12,600 units × $30 per unit) $ 378,000 Variable expenses 226,800 Contribution margin 151,200 Fixed expenses 169,200 Net operating loss $ (18,000 )

Required: 1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales. 2. The president believes that a $6,400 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $81,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income? 3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $30,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)? 4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.60 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,200? 5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $54,000 each month. a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. b. Assume that the company expects to sell 20,500 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,500)?
Answer #1

ANSWER

Solution 1:

CM ratio = Contribution margin / Sales = $151,200 / $378,000 = 40%

break-even point in unit sales = Fixed expenses / CM per unit = $169,200 / $12 = 14100 units

Breakeven point in dollar sales = Fixed expenses / CM ratio = $169,200 / 40% = $423,000

Solution 2:

increase (decrease) in the company’s monthly net operating income = Increase in contribution margin - Increase in advertising expenses

= ($81,000*40%) - $6,400 = $26,000

Solution 3:

New selling price per unit = $30*90% = $27

New CM per unit = $27 - $18 = $9 per unit

New fixed costs =$169,200 + $30,000 = $199,200

New sales volume = 12600*2 = 25200 units

New operating income = Contribution margin - Fixed costs = (25200*$9) - $199,200 = $27,600

Solution 4:

New contribution margin per unit = $12 - $0.60 = $11.40 per unit

Target income = $4,200

Nos of units to be sold to attain target profit = (Fixed cost + Net operating income) / CM per unit

= ($169,200 + $4,200) / $11.40 = 15,210 units

-----------------------------------------

Note: I have answered first 4 parts of the question as HOMEWORKLIB RULES policy, kindly post separate question for answer of remaining parts.

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************